I need to know the formula for adding whole numbers (1+2+3+4...) to get a sum over 1000. I know the answer is 45 but I need the formula to figure it out.


Answer by ad_alta(240)

I should probably let you discover it on your own (its more fun that way), but the answer is n(n+1)/2. Since we're talking about n=1000, 1000(1001)/2=500,500.
Or in your case, (1+2+3+4+5+6+7+8+9)=(9)(10)/2=45.